A grocer wishes to pile cans of ice cream so that the top layer will contain three cans and each layer will have one more can than the layer above. If there are 10 layers, how many cans would he need ?

Answer by amarjeeth123(569)

The given number of cans=3+4+.........+12
=1+2+3+......+12-(1+2)
=12(12+1)/2-3
=6(13)-3
=78-3
=75
The grocer would need 75 cans.